# A mathematical theory of topological invariants of quantum spin systems

### v2

@@ 5.2 Symmetries of lattice systems
If $\Lambda\subset{\mathbb{R}}^{n}$ is countable, it can be viewed as a lattice in the physical sense, and the local Lie algebra ${{\mathfrak{g}}_{al}}$ over ${\mathcal{C}\mathcal{S}}_{n}$ models infinitesimal gauge transformations of a lattice system on ${\mathbb{R}}^{n}$.
Definition 5.1.
@@
A local action of a compact Lie group on a lattice system $(\Lambda,\{V_{j}\}_{j\in\Lambda})$ is a collection of homomorphisms $\rho_{j}:G\rightarrow U(V_{j})$ such that the norms of the corresponding Lie algebra homomorphisms ${\mathfrak{g}}\rightarrow B(V_{j})$ are bounded uniformly in $j$.
@@
A local action of $G$ on $(\Lambda,\{V_{j}\}_{j\in\Lambda})$ gives rise to a homomorphism from $G$ to the automorphism group of ${\mathscr{A}}$ via $g\mapsto\otimes_{j\in\Lambda}{\rm Ad}_{\rho_{j}(g)}$ which is smooth on ${\mathscr{A}}_{a\ell}$. The corresponding generator ${\mathsf{Q}}$ is a homomorphism from ${\mathfrak{g}}$ to the Lie algebra of derivations of ${\mathscr{A}}_{a\ell}$ defined by
${\mathsf{Q}}:(a,{\mathcal{A}})\mapsto\sum_{j\in\Lambda}[{\mathsf{q}}_{j}(a),{\mathcal{A}}],\quad a\in{\mathfrak{g}},\ {\mathcal{A}}\in{\mathscr{A}}_{a\ell},$
where ${\mathsf{q}}_{j}$ is the traceless part of the generator of $\rho_{j}$. The image of ${\mathsf{Q}}$ lands in ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$, with ${\mathsf{Q}}(a)^{Y}=\sum_{j\in\Lambda}{\mathsf{q}}_{j}(a)^{Y}$. We will regard ${\mathsf{Q}}$ as a homomorphism of Fréchet-Lie algebras ${\mathfrak{g}}\rightarrow{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$.
@@
This morphism of Fréchet-Lie algebras can be lifted to a morphism of local Lie algebras ${{\mathfrak{g}}_{al}}\rightarrow{\mathfrak{D}}_{al}$ over ${\mathcal{C}\mathcal{S}}_{n}$. Indeed, for any $U\in{\mathcal{C}\mathcal{S}}_{n}$ and any $f\in{{\mathfrak{g}}_{al}}(U)$ we let ${\mathsf{Q}}(f)$ be a derivation of ${\mathscr{A}}_{a\ell}$ given by
${\mathsf{Q}}(f)({\mathcal{A}})=\sum_{j\in\Lambda}[{\mathsf{q}}_{j}(f(j)),{\mathcal{A}}],\quad{\mathcal{A}}\in{\mathscr{A}}_{a\ell}.$
@@
It is easy to check that this derivation belongs to ${\mathfrak{D}}_{al}(U)$ and that the above map is a continuous homomorphism ${{\mathfrak{g}}_{al}}(U)\rightarrow{\mathfrak{D}}_{al}(U)$. The physical interpretation is that a local action of a compact Lie group on a quantum lattice system can be gauged on the infinitesimal level.
Definition 5.2.
@@
A state $\psi$ of ${\mathscr{A}}$ is said to be invariant under a local action of a compact Lie group $G$ if it is invariant under the corresponding automorphisms of ${\mathscr{A}}$.
@@
Let $\psi$ be a gapped state of ${\mathscr{A}}$ invariant under a local action of a Lie group $G$. In that case the image of ${\mathsf{Q}}:{\mathfrak{g}}\rightarrow{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ lands in ${\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})$. One may ask if this morphism of Fréchet-Lie algebras can be lifted to a morphism of local Lie algebras ${{\mathfrak{g}}_{al}}\rightarrow{\mathfrak{D}}^{\psi}_{al}$. If this is the case, then the symmetry $G$ of $\psi$ can be gauged on the infinitesimal level. In the next section we construct obstructions for the existence of such a morphism of local Lie algebras and show that zero-temperature Hall conductance is an example of such an obstruction.
@@ 5.3 Equivariantization
As a preliminary step, for any $G$-invariant gapped state $\psi$ we are going to define a graded local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$ which is a $G$-equivariant version of the local Lie algebra ${\mathfrak{D}}^{\psi}_{al}$. Recall that a graded local Lie algebra is a cosheaf of graded vector spaces that is a pre-cosheaf of graded Lie algebras satisfying the graded analogue of Property I. For example, if ${\mathfrak{F}}$ is a local Lie algebra and $A=\prod_{k\in{\mathbb{Z}}}A_{k}$ is a locally finite supercommutative graded algebra with finite-dimensional graded factors $A_{k}$, [^10] A graded vector space is locally finite iff its graded components are finite-dimensional. then $U\mapsto{\mathfrak{F}}(U)\otimes A$ is a graded local Lie algebra. We denote it ${\mathfrak{F}}\otimes A$.
Fix a compact Lie group $G$ and a distributive pre-ordered set $X$ and consider the category of graded local Lie algebras over $X$ equipped with a $G$-action. An object of this category is a graded local Lie algebra ${\mathfrak{F}}$ on which $G$ acts by automorphisms; morphisms are defined in an obvious manner. The first step is to define a functor ${\mathfrak{F}}\mapsto{\mathfrak{F}}^{G}$ from this category to the category of graded local Lie algebras over $X$ such that ${\mathfrak{F}}^{G}(U)$ is the Lie algebra of $G$-invariant elements of ${\mathfrak{F}}(U)$. It is clear how to define such a functor for coflasque pre-cosheaves of Lie algebras with Property I, but the pre-cosheaf ${\mathfrak{F}}^{G}$ will not be a cosheaf of vector spaces without further assumptions about ${\mathfrak{F}}$ and the $G$-action.
Definition 5.3.
An action of $G$ on a pre-cosheaf of Fréchet spaces ${\mathfrak{F}}$ is smooth if for each $U\in X$ the seminorms defining the topology of ${\mathfrak{F}}(U)$ can be chosen to be $G$-invariant and the map $G\times{\mathfrak{F}}(U)\rightarrow{\mathfrak{F}}(U)$ defining the action is smooth. An action of $G$ on a pre-cosheaf of graded Fréchet spaces is smooth is the $G$-action on every graded component is smooth.
Proposition 5.2.
Let ${\mathfrak{F}}$ be a pre-cosheaf of graded Fréchet  spaces over $X$ equipped with a smooth action of a compact Lie group $G$. The assignment $U\mapsto{\mathfrak{F}}^{G}(U)=\left({\mathfrak{F}}(U)\right)^{G}$ is a cosheaf of graded Fréchet spaces.
@@
It is sufficient to prove this in the ungraded case. We need to show that for any $U,V\in X$ the sequence
[…]
is exact. To show exactness at the rightmost term, let ${\mathsf{F}}$ be a $G$-invariant element of ${\mathfrak{F}}(U\vee V)$ and let ${\mathsf{F}}_{U}\in{\mathfrak{F}}(U)$ and ${\mathsf{F}}_{V}\in{\mathfrak{F}}(V)$ be such that $\iota_{U\cup V,U}{\mathsf{F}}_{U}+\iota_{U\cup V,V}{\mathsf{F}}_{V}={\mathsf{F}}$. Averaging the action map $G\times{\mathfrak{F}}(U)$ over $G$ with the Haar measure gives a linear map $h_{U}:{\mathfrak{F}}(U)\mapsto{\mathfrak{F}}^{G}(U)$ which is identity when restricted to ${\mathfrak{F}}^{G}(U)$. The co-restriction morphisms intertwine these maps. Thus $\iota_{U\cup V,U}\circ h_{U}({\mathsf{F}}_{U})+\iota_{U\cup V,V}\circ h_{V}({\mathsf{F}}_{V})={\mathsf{F}}$ which proves that $\beta$ is surjective. Exactness in the middle term is proved similarly. ∎
Remark 5.1.
For the proof to go through, it suffices to require the map $G\times{\mathfrak{F}}(U)\rightarrow{\mathfrak{F}}(U)$ to be continuous. However, if it is smooth, ${\mathfrak{F}}(U)$ becomes a ${\mathfrak{g}}$-module and all elements in ${\mathfrak{F}}^{G}(U)\subset{\mathfrak{F}}(U)$ are annihilated by the ${\mathfrak{g}}$-action. We will use this later on.
Example 5.1.
If $G$ acts locally on a lattice system, the action of $G$ on the local Lie algebra ${\mathfrak{D}}_{al}$ is smooth. If $\psi$ is a $G$-invariant gapped state of such a lattice system, the action of $G$ on ${\mathfrak{D}}^{\psi}_{al}$ is smooth.
Example 5.2.
Consider the local Lie algebra ${{\mathfrak{g}}_{al}}$ over ${\mathcal{C}\mathcal{S}}_{n}$ associated to a finite-dimensional Lie algebra ${\mathfrak{g}}$ (see Section 4). Assume that ${\mathfrak{g}}$ is the Lie algebra of a compact Lie group $G$, then there is an obvious $G$-action on ${{\mathfrak{g}}_{al}}$:
$(g\cdot f)(j)={\rm Ad}_{g}f(j),\quad g\in G,j\in\Lambda.$
@@
If ${\mathfrak{F}}$ is a local Fréchet-Lie algebra with a smooth $G$-action and $A=\prod_{k\in{\mathbb{Z}}}A_{k}$ is a locally-finite supercommutative graded algebra on which $G$ acts by automorphisms, then the $G$-action on ${\mathfrak{F}}\otimes A$ is smooth.
Corollary 5.1.
@@
Let ${\mathfrak{F}}$ be a graded local Fréchet-Lie algebra over $X$ with a smooth $G$-action. The functor of $G$-invariant elements maps ${\mathfrak{F}}$ to a graded local FréchetLie algebra ${\mathfrak{F}}^{G}$ over $X$.
Definition 5.4.
@@
Let ${\mathfrak{F}}$ be a local Fréchet-Lie algebra over $X$ with a smooth $G$-action. The $G$-equivariantization functor sends ${\mathfrak{F}}$ to the negatively-graded local Fréchet-Lie algebra ${\mathfrak{F}}^{\bf G}$ defined by
$U\mapsto\left({\mathfrak{F}}(U)\otimes\prod_{k=1}^{\infty}{\rm Sym}^{k}({\mathfrak{g}}^{*}[-2])\right)^{G}.$
@@
In the cases of interest to us, the $G$-action on a local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$ is infinitesimally inner, in the sense that the ${\mathfrak{g}}$-module structure mentioned in Remark 5.1 arises from a homomorphism $\rho:{\mathfrak{g}}\rightarrow{\mathfrak{F}}({\mathbb{R}}^{n})$. In such a case, the graded local Lie algebra ${\mathfrak{F}}^{\bf G}$ has an extra bit of structure: a central element in ${\mathfrak{F}}^{\bf G}({\mathbb{R}}^{n})$ of degree $-2$. This element is simply $\rho$ re-interpreted as an element of ${\mathfrak{F}}({\mathbb{R}}^{n})\otimes{\mathfrak{g}}^{*}[-2]$. In the terminology of Section 2, ${\mathfrak{F}}^{\bf G}$ is a pointed graded local Fréchet-Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$. It is easy to see that the $G$-equivariantization functor respects this extra structure. That is, if $f:{\mathfrak{F}}\rightarrow{\mathfrak{F}}^{\prime}$ is a morphism of local Fréchet-Lie algebras over ${\mathcal{C}\mathcal{S}}_{n}$ commuting with infinitesimally inner smooth $G$-actions on ${\mathfrak{F}}$ and ${\mathfrak{F}}^{\prime}$, then $f^{\bf G}$ maps the central element $\rho\in{\mathfrak{F}}^{\bf G}({\mathbb{R}}^{n})_{-2}$ to the central element $\rho^{\prime}\in{{\mathfrak{F}}^{\prime}}^{\bf G}({\mathbb{R}}^{n})_{-2}$.
Example 5.4.
@@
Let $\psi$ be a $G$-invariant gapped state of a quantum lattice system with a local $G$-action which on the infinitesimal level is described by ${\mathsf{Q}}:{\mathfrak{g}}\rightarrow{\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})$. Consider the graded local Lie algebra ${\mathfrak{D}}^{\psi}_{al}$ with its smooth $G$-action (Example 5.1) and its $G$-equivariantization ${\mathfrak{D}}^{\psi,\bf G}_{al}$. The distinguished central element of ${\mathfrak{D}}^{\psi,\bf G}_{al}({\mathbb{R}}^{n})$ is ${\mathsf{Q}}$ regarded as an element of ${\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})\otimes{\mathfrak{g}}^{*}[-2]$.
Example 5.5.
@@
Consider the graded local Lie algebra ${{\mathfrak{g}}^{\bf G}_{al}}$ of Example 5.2. The degree $-2$ component of ${{\mathfrak{g}}^{\bf G}_{al}}({\mathbb{R}}^{n})$ is the space of $G$-invariant bounded functions on $\Lambda$ with values in ${\mathfrak{g}}\otimes{\mathfrak{g}}^{*}$. The distinguished central element is the constant function on $\Lambda$ which takes the value ${\rm id}_{\mathfrak{g}}$.
@@
Armed with the equivariantization functor, we can now explain our strategy for constructing obstructions for the existence of a local Lie algebra morphism ${{\mathfrak{g}}_{al}}\rightarrow{\mathfrak{D}}^{\psi}_{al}$ which lifts the Fréchet-Lie algebra morphism ${\mathsf{Q}}:{\mathfrak{g}}\rightarrow{\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})$. Suppose such a morphism $\rho$ exists. Applying the $G$-equivariantization functor, we get a morphism of pointed negatively-graded local Fréchet-Lie algebras $\rho^{\bf G}:{{\mathfrak{g}}^{\bf G}_{al}}\rightarrow{\mathfrak{D}}^{\psi,\bf G}_{al}$. For any CS cover ${\mathfrak{U}}$ of ${\mathbb{R}}^{n}$ an application of the Čech functor gives a morphism of acyclic pointed DGLAs
$C^{aug}_{\bullet+1}({\mathfrak{U}},{\mathbb{R}}^{n};{{\mathfrak{g}}^{\bf G}_{al}})\rightarrow C^{aug}_{\bullet+1}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,\bf G}_{al}).$
@@
Consequently, a obstruction for the existence of such a pointed DGLA morphism is an obstruction for the existence of $\rho$. In the next section we use the twisted Maurer-Cartan equation for pointed DGLAs to construct such obstructions and identify them as topological invariants of gapped states defined in [11].

### v3

@@ 5.2 Symmetries of lattice systems
If $\Lambda\subseteq{\mathbb{R}}^{n}$ is countable, it can be viewed as a lattice in the physical sense, and the local Lie system ${\mathfrak{g}}_{al}$} over ${\mathcal{C}\mathcal{S}}_{n}$ models infinitesimal gauge transformations of a lattice system on ${\mathbb{R}}^{n}$. First, we generalize Example 5 by replacing $U(1)$ with an arbitrary compact Lie group.
Definition 90.
@@
An on-site action of a compact Lie group $G$ on a lattice system $(\Lambda,\{V_{x}\}_{x\in\Lambda})$ is a collection of continuous (and therefore smooth) homomorphisms $\rho_{x}:G\rightarrow U(V_{x})^{even}$ such that the norms of the corresponding Lie algebra homomorphisms ${\mathfrak{g}}\rightarrow B(V_{x})$ are bounded uniformly in $x$.
@@
An on-site action of $G$ on $(\Lambda,\{V_{x}\}_{x\in\Lambda})$ gives rise to a homomorphism from $G$ to the automorphism group of $\mathscr{A}$$_{\ell}$ via $g\mapsto\otimes_{x\in\Lambda}{\rm Ad}_{\rho_{x}(g)}$.
Lemma 91.
An on-site action of $G$ on $\mathscr{A}$$_{\ell}$ extends to an action on ${\mathscr{A}}_{a\ell}$. This action is smooth in the sense that the corresponding map $G\times{\mathscr{A}}_{a\ell}\rightarrow{\mathscr{A}}_{a\ell}$ is smooth with respect to the Fréchetnorm. The generator of the action is a homomorphism ${\mathfrak{g}}\rightarrow{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ which sends $a\in{\mathfrak{g}}$ to a function o
$Y\mapsto\sum_{x\in Y\cap\Lambda}{\mathsf{q}}_{x}(a)^{Y}.$ (119)
where $Y\in{\mathbb{B}}_{n}$ and ${\mathsf{q}}_{x}$ is the traceless part of the generator of $\rho_{x}$.
Proof.
Since ${\mathscr{A}}_{a\ell}$ is the completion of $\mathscr{A}$$_{\ell}$ with respect to the norms (70) and any on-site $G$ action preserves these norms, the action clearly extends to ${\mathscr{A}}_{a\ell}$. Let us denote by $\rho$ the corresponding map from $G$ to the space of linear maps $L({\mathscr{A}}_{a\ell},{\mathscr{A}}_{a\ell})$. To prove the second statement, note that the map $\rho$ is continuous in the strong operator topology. Then it is a standard fact that to prove joint smoothness, it is enough to prove smoothness of the function $g\mapsto\rho(g)({\mathcal{A}})$ for a fixed ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}$ [22]. The latter is straightforward, since the derivatives are given by explicit formulas. For example, the 1st derivative is
$d\rho(g)({\mathcal{A}})(a)=\rho(g)({\mathsf{Q}}(a)({\mathcal{A}})),\quad g\in G,a\in{\mathfrak{g}},{\mathcal{A}}\in{\mathscr{A}}_{a\ell},$ (120)
where
${\mathsf{Q}}(a)({\mathcal{A}})=\sum_{x\in\Lambda}[{\mathsf{q}}_{x}(a),{\mathcal{A}}].$ (121)
Since $\rho(g)$ is continuous in strong operator topology, so is its first derivative. The continuity of higher derivatives is proved similarly. Setting $g$ to be the identity of $G$, we find that the generator of the action is given by $\mathsf{Q}$}, and it is easy to see that the corresponding function ${\mathbb{B}}_{n}\rightarrow{{\mathfrak{d}}_{l}}$ is given by (119). ∎
@@
This morphism of Fréchet-Lie algebras can be lifted to a morphism of local Lie systems ${{\mathfrak{g}}_{al}}\rightarrow{\mathfrak{D}}_{al}$ over ${\mathcal{C}\mathcal{S}}_{n}$. Indeed, for any $U\in{\mathcal{C}\mathcal{S}}_{n}$ and any $f\in{{\mathfrak{g}}_{al}}(U)$ we let ${\mathsf{Q}}(f)$ be a derivation of ${\mathscr{A}}_{a\ell}$ given by
${\mathsf{Q}}(f)({\mathcal{A}})=\sum_{x\in\Lambda}[{\mathsf{q}}_{x}(f(x)),{\mathcal{A}}],\quad{\mathcal{A}}\in{\mathscr{A}}_{a\ell}.$ (122)
@@
It is easy to check that this derivation belongs to ${\mathfrak{D}}_{al}(U)$ and that the above map is a continuous homomorphism ${{\mathfrak{g}}_{al}}(U)\rightarrow{\mathfrak{D}}_{al}(U)$. The physical interpretation is that an on-site action of a compact Lie group on a quantum lattice system can be gauged on the infinitesimal level.
Definition 92.
@@
A state $\psi$ of $\mathscr{A}$} is said to be invariant under an on-site action of a compact Lie group $G$ if it is invariant under the corresponding automorphisms of $\mathscr{A}$}.
@@
Let $\psi$ be a gapped state of $\mathscr{A}$} invariant under an on-site action of a compact Lie group $G$. In that case the image of ${\mathsf{Q}}:{\mathfrak{g}}\rightarrow{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ lands in ${\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})$. One may ask if this morphism of Fréchet-Lie algebras can be lifted to a morphism of local Lie systems ${{\mathfrak{g}}_{al}}\rightarrow{\mathfrak{D}}^{\psi}_{al}$. If this is the case, then the symmetry $G$ of $\psi$ can be gauged on the infinitesimal level. In the next section we construct obstructions for the existence of such a morphism of local Lie systems and show that zero-temperature Hall conductance is an example of such an obstruction.
@@ 5.3 Equivariantization
As a preliminary step, for any $G$-invariant gapped state $\psi$ we are going to define a graded local Lie system over ${\mathcal{C}\mathcal{S}}_{n}$ which is a $G$-equivariant version of the local Lie system ${\mathfrak{D}}^{\psi}_{al}$. Recall that a graded local Lie system is a cosheaf of graded vector spaces that is a pre-cosheaf of graded Lie algebras satisfying the graded analogue of Property I. For example, if $\mathfrak{F}$} is a local Lie system and $A=\prod_{k\in{\mathbb{Z}}}A_{k}$ is a locally finite supercommutative graded algebra with finite-dimensional graded factors $A_{k}$, [^12] A graded vector space is locally finite iff its graded components are finite-dimensional. then $U\mapsto{\mathfrak{F}}(U)\otimes A$ is a graded local Lie system. We denote it ${\mathfrak{F}}\otimes A$.
Fix a compact Lie group $G$ and a distributive pre-ordered set $X$ and consider the category of graded local Lie systems over $X$ equipped with a $G$-action. An object of this category is a graded local Lie system $\mathfrak{F}$} on which $G$ acts by automorphisms; morphisms are defined in an obvious manner. The first step is to define a functor ${\mathfrak{F}}\mapsto{\mathfrak{F}}^{G}$ from this category to the category of graded local Lie systems over $X$ such that ${\mathfrak{F}}^{G}(U)$ is the Lie algebra of $G$-invariant elements of ${\mathfrak{F}}(U)$. It is clear how to define such a functor for coflasque pre-cosheaves of Lie algebras with Property I, but the pre-cosheaf ${\mathfrak{F}}^{G}$ will not be a cosheaf of vector spaces without further assumptions about $\mathfrak{F}$} and the $G$-action.
Definition 93.
An action of $G$ on a pre-cosheaf of Fréchet spaces $\mathfrak{F}$} is smooth if for each $U\in X$ the seminorms defining the topology of ${\mathfrak{F}}(U)$ can be chosen to be $G$-invariant and the map $G\times{\mathfrak{F}}(U)\rightarrow{\mathfrak{F}}(U)$ defining the action is smooth. An action of $G$ on a pre-cosheaf of graded Fréchet spaces is smooth if the $G$-action on every graded component is smooth.
Proposition 94.
Let $\mathfrak{F}$} be a cosheaf of graded Fréchet spaces over $X$ equipped with a smooth action of a compact Lie group $G$. The assignment $U\mapsto{\mathfrak{F}}^{G}(U)=\left({\mathfrak{F}}(U)\right)^{G}$ is a cosheaf of graded Fréchet spaces.
@@
We need to show that for any $U,V\in X$ the sequence
${\mathfrak{F}}^{G}(U\wedge V)\xrightarrow{\alpha}{\mathfrak{F}}^{G}(U)\oplus{\mathfrak{F}}^{G}(V)\xrightarrow{\beta}{\mathfrak{F}}^{G}(U\vee V)\to 0,$ (123)
is exact. To show exactness at the rightmost term, let $\mathsf{F}$} be a $G$-invariant element of ${\mathfrak{F}}(U\vee V)$ and let ${\mathsf{F}}_{U}\in{\mathfrak{F}}(U)$ and ${\mathsf{F}}_{V}\in{\mathfrak{F}}(V)$ be such that $\iota_{U\vee V,U}{\mathsf{F}}_{U}+\iota_{U\vee V,V}{\mathsf{F}}_{V}={\mathsf{F}}$. Averaging the action map $G\times{\mathfrak{F}}(U)$ over $G$ with the Haar measure gives a linear map $h_{U}:{\mathfrak{F}}(U)\rightarrow{\mathfrak{F}}^{G}(U)$ which is identity when restricted to ${\mathfrak{F}}^{G}(U)$. The co-restriction morphisms intertwine these maps. Thus $\iota_{U\vee V,U}\circ h_{U}({\mathsf{F}}_{U})+\iota_{U\vee V,V}\circ h_{V}({\mathsf{F}}_{V})={\mathsf{F}}$ which proves that $\beta$ is surjective. Exactness in the middle term is proved similarly. ∎
Remark 95.
For the proof to go through, it suffices to require the map $G\times{\mathfrak{F}}(U)\rightarrow{\mathfrak{F}}(U)$ to be continuous. However, if it is smooth, ${\mathfrak{F}}(U)$ becomes a $\mathfrak{g}$}-module and all elements in ${\mathfrak{F}}^{G}(U)\subseteq{\mathfrak{F}}(U)$ are annihilated by the $\mathfrak{g}$}-action. We will use this later on.
Example 96.
If $G$ acts on-site on a lattice system, the action of $G$ on the local Lie system ${\mathfrak{D}}_{al}$ is smooth. If $\psi$ is a $G$-invariant gapped state of such a lattice system, the action of $G$ on ${\mathfrak{D}}^{\psi}_{al}$ is smooth.
Example 97.
Consider the local Lie system ${\mathfrak{g}}_{al}$} over ${\mathcal{C}\mathcal{S}}_{n}$ associated to a finite-dimensional Lie algebra $\mathfrak{g}$} (see Section 4). Assume that $\mathfrak{g}$} is the Lie algebra of a compact Lie group $G$, then there is an obvious $G$-action on ${\mathfrak{g}}_{al}$}:
$(g\cdot f)(x)={\rm Ad}_{g}f(x),\quad g\in G,x\in\Lambda.$ (124)
@@
If $\mathfrak{F}$} is a local Fréchet-Lie system with a smooth $G$-action and $A=\prod_{k\in{\mathbb{Z}}}A_{k}$ is a locally-finite supercommutative graded algebra on which $G$ acts by automorphisms, then the $G$-action on ${\mathfrak{F}}\otimes A$ is smooth.
Corollary 99.
@@
Let $\mathfrak{F}$} be a graded local Fréchet-Lie system over $X$ with a smooth $G$-action. The functor of $G$-invariant elements maps $\mathfrak{F}$} to a graded local FréchetLie system ${\mathfrak{F}}^{G}$ over $X$.
Definition 100.
@@
Let $\mathfrak{F}$} be a local Fréchet-Lie system over $X$ with a smooth $G$-action. The $G$-equivariantization functor sends $\mathfrak{F}$} to the negatively-graded local Fréchet-Lie system ${\mathfrak{F}}^{\bf G}$ defined by
$U\mapsto\left({\mathfrak{F}}(U)\otimes\prod_{k=1}^{\infty}{\rm Sym}^{k}({\mathfrak{g}}^{*}[-2])\right)^{G}.$ (125)
@@
In the cases of interest to us, the $G$-action on a local Lie system over ${\mathcal{C}\mathcal{S}}_{n}$ is infinitesimally inner, in the sense that the $\mathfrak{g}$}-module structure mentioned in Remark 95 arises from a homomorphism $\rho:{\mathfrak{g}}\rightarrow{\mathfrak{F}}({\mathbb{R}}^{n})$. In such a case, the graded local Lie system ${\mathfrak{F}}^{\bf G}$ has an extra bit of structure: a central element in ${\mathfrak{F}}^{\bf G}({\mathbb{R}}^{n})$ of degree 2-2. This element is simply $\rho$ re-interpreted as an element of ${\mathfrak{F}}({\mathbb{R}}^{n})\otimes{\mathfrak{g}}^{*}[-2]$. In the terminology of Section 3, ${\mathfrak{F}}^{\bf G}$ is a pointed graded local Fréchet-Lie system over ${\mathcal{C}\mathcal{S}}_{n}$. It is easy to see that the $G$-equivariantization functor respects this extra structure. That is, if $f:{\mathfrak{F}}\rightarrow{\mathfrak{F}}^{\prime}$ is a morphism of local Fréchet-Lie system over ${\mathcal{C}\mathcal{S}}_{n}$ commuting with infinitesimally inner smooth $G$-actions on $\mathfrak{F}$} and $\mathfrak{F}$$^{\prime}$, then $f^{\bf G}$ maps the central element $\rho\in{\mathfrak{F}}^{\bf G}({\mathbb{R}}^{n})_{-2}$ to the central element $\rho^{\prime}\in{{\mathfrak{F}}^{\prime}}^{\bf G}({\mathbb{R}}^{n})_{-2}$.
Example 101.
@@
Let $\psi$ be a $G$-invariant gapped state of a quantum lattice system with an on-site $G$-action which on the infinitesimal level is described by ${\mathsf{Q}}:{\mathfrak{g}}\rightarrow{\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})$. Consider the local Lie system ${\mathfrak{D}}^{\psi}_{al}$ with its smooth $G$-action (Example 96) and its $G$-equivariantization ${\mathfrak{D}}^{\psi,\bf G}_{al}$. The distinguished central element of ${\mathfrak{D}}^{\psi,\bf G}_{al}({\mathbb{R}}^{n})$ is $\mathsf{Q}$} regarded as an element of ${\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})\otimes{\mathfrak{g}}^{*}[-2]$.
Example 102.
@@
Consider the local Lie system ${\mathfrak{g}}_{al}$} of Example 97. The degree 2-2 component of ${{\mathfrak{g}}^{\bf G}_{al}}({\mathbb{R}}^{n})$ is the space of $G$-invariant bounded functions on $\Lambda$ with values in ${\mathfrak{g}}\otimes{\mathfrak{g}}^{*}$. The distinguished central element is the constant function on $\Lambda$ which takes the value ${\rm id}_{\mathfrak{g}}$.
@@
Armed with the equivariantization functor, we can now explain our strategy for constructing obstructions for the existence of a local Lie system morphism ${{\mathfrak{g}}_{al}}\rightarrow{\mathfrak{D}}^{\psi}_{al}$ which lifts the Fréchet-Lie algebra morphism ${\mathsf{Q}}:{\mathfrak{g}}\rightarrow{\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})$. Suppose such a morphism $\rho$ exists. Applying the $G$-equivariantization functor, we get a morphism of pointed negatively-graded local Fréchet-Lie systems $\rho^{\bf G}:{{\mathfrak{g}}^{\bf G}_{al}}\rightarrow{\mathfrak{D}}^{\psi,\bf G}_{al}$. For any ${\mathcal{C}\mathcal{S}}_{n}$-cover $\mathfrak{U}$} of ${\mathbb{R}}^{n}$ an application of the Čech functor gives a morphism of acyclic pointed DGLAs
$C^{aug}_{\bullet+1}({\mathfrak{U}},{\mathbb{R}}^{n};{{\mathfrak{g}}^{\bf G}_{al}})\rightarrow C^{aug}_{\bullet+1}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,\bf G}_{al}).$ (126)
@@
Consequently, an obstruction for the existence of such a pointed DGLA morphism is an obstruction for the existence of $\rho$. In the next section we use the twisted Maurer-Cartan equation for pointed DGLAs to construct such obstructions and identify them as topological invariants of gapped states defined in [19].